OpenStudy (anonymous):

I am having a lot of trouble rearranging an equation to find the angle needed to reach a certain horizontal range. I must use equation R=v(squared)/g times (sin 2 theta). I know the max range is 700 meters, and that sin theta can be no greater than 1. Please help!

OpenStudy (anonymous):

\[\frac{ d*g }{ V ^{2} } = \sin 2\theta \] \[\sin^{-1} \frac{ d*g }{ V ^{2} } = 2\theta \] \[\frac{ \sin^{-1}( \frac{ dg }{ V ^{2} } ) }{ 2 } =\theta \]

OpenStudy (anonymous):

Thank you, I kept trying to put a square root in! You are wonderful